# Distributed Maximum Flow in Planar Graphs

### v3

@@ 5.1.3 Distributed Knowledge Properties
In $\tilde{O}(D)$ rounds, for each bag $X$, all vertices know for each of their incident copies of edges whether they participate in the critical face (face-part) and whether they participate in any face-part of $X$.
@@
After each vertex $v$ knows the necessary information for each copy of an incident edge $e$, in an additional round, $v$ knows if $e$ exists in $X^{*}$, its weight in $X^{*}$ and its direction. Proving property 14 of the BDD.
@@ 5.2 The Labeling Scheme
Here we describe our distance labeling scheme. Recall that, for every dual bag $X^{*}$, we wish to label the nodes of $X^{*}$, such that, from the labels alone of any two nodes we can deduce their distance in $X^{*}$. Recall, we refer to the dual node $f^{\prime}$ in a child bag $X^{*}_{i}$ as a node-part of a node $f\in X^{*}$, if $f^{\prime}$ corresponds to a face-part in $X_{i}$ resulting from partitioning the face or face-part $f$ of $X$.
We already know that this set would play the role of a node-cut (separator) of $X^{*}$, and is crucial for the labeling scheme and algorithm. The main structural property that $F_{\!{}_{X}}$ provides is: Any path in $X^{*}$ either intersects $F_{\!{}_{X}}$ or is entirely contained in a child bag of $X^{*}$.
Note, a node $g\notin F_{\!{}_{X}}$ of $X^{*}$ corresponds to real faces of $G$ (i.e. not face-parts) contained in $X$. The distance label $\mathsf{Label}_{X^{*}}(g)$ of $g\notin F_{\!{}_{X}}$ is defined recursively. If $X^{*}$ is a leaf-bag then $\mathsf{Label}_{X^{*}}(g)$ stores ID($g$) and the distances between $g$ and all other nodes $h\in X^{*}$. Otherwise, the label consists of the ID of $X$, the distances in $X^{*}$ between $g$ and all nodes of $F_{\!{}_{X}}$, and (recursively) the label of $g$ in the unique (by Corollary 5.6) child bag $X^{*}_{i}$ of $X^{*}$ that entirely contains $g$. Formally:
@@
[…]
@@
The label of any node $g$ in a bag $X^{*}$ of $X\in\mathcal{T}$ is of size $\tilde{O}(D)$.
@@ 5.3 The Labeling Algorithm
3forall bags $X$ of the same level, bottom-up, in parallel,  do
@@ 11            Broadcast the dual edges of $S_{\!{}_{X}}$ edges that exist in $X^{*}$ to the entire bag $X$;
19      Using the previously received information, each vertex $v\in X$ computes locally $\mathsf{Label}_{X^{*}}(g)$ for each face (face-part) $g$ of $X$ that contains it or reports a negative cycle aborting the algorithm;
@@
[…]
@@
  • A bidirectional zero weighted edge between each two nodes $f_{i}\in V_{i}$, $f_{j}\in V_{j}$ that correspond to node-parts in $X_{i}$, $X_{j}$ (respectively) of the same node $f$ in $F_{\!{}_{X}}$. Again, this can be known locally in any vertex $v$ of $X$ since $v$ received the labels (which contain the IDs) of $f_{i},f_{j}$ in the broadcast step, and since the correspondence of $f_{i},f_{j}$ to $f$ can be known from the information received in the broadcast (specifically, the IDs of $f_{i},f_{j}$ contain the ID of $f$. For more details, see the proof of Lemma 5.12 in Appendix C).
@@
To conclude, a BDD $\mathcal{T}$ is built in $\tilde{O}(D)$ rounds with each vertex $v\in G$ knowing all relevant information that is locally needed about $X^{*}$ (i.e., $X^{*}$ is distributively stored). In each level of $\mathcal{T}$ we have at most $\tilde{O}(D^{2})$ rounds. Thus, by Lemmas 5.18 and 5.19 the above gives a correct distance labeling algorithm of the dual network. Formally,
@@ 5.3.1 SSSP in the Dual
To mark the SSSP tree edges, an aggregate operator is computed for each node $v$ of $G^{*}$ over its neighbors (using Lemma B.9, we perform a single part-wise aggregation task). In particular, for a dual node, we mark its incident edge that minimizes the distance product with its neighbors, i.e, we take the operator computed is $\min_{u\in N(v)}\{\mathsf{dist}(s,u)+w(u,v)\}$ (where $N(v)$ is $v$’s neighborhood). Finally, $G^{*}$ might be a multi-graph, so there might be many edges connecting $v$ with a neighbor. We mark the minimal weight edge connecting $v$ to the neighbor that minimizes the above (breaking ties by IDs for edges and for neighbors). Thus,
@@
The above theorem in addition to Section 4 imply distributed (exact and approximate) algorithms for Maximum $st$-Flow in $G$ (i.e., proves Theorems 1.2 and 1.3) by an adaptation of the centralized algorithms of Miller-Naor [30] and of Hassin [20]. The details are given in Section 6.
@@ 6.1 Maximum st-Flow
The proof mostly follows a centralized algorithm of Miller and Naor [30], we provide an overview of the general idea. For the full proof see Section D.1. The algorithm uses a logarithmic number of SSSP computations in the dual, after which one can deduce the value of the max $st$-flow and a flow assignment. In particular, it performs a binary search on the value $\lambda$ of the maximum $st$-flow. In each iteration, $\lambda$ units of flow are pushed along a (not necessarily shortest nor directed) $s$-to-$t$ path. If this does not violate the capacity of any dart in $P$ then it is a feasible flow and we conclude that the maximum $st$-flow is at least $\lambda$. Otherwise, a residual graph is defined, possibly having negative capacities. To fix that, an SSSP computation from an arbitrary node $f$ in the dual is performed, such that, the flow can be fixed to not violate capacities if and only if the dual does not contain a negative cycle (if a negative cycle was found, then, the max $st$-flow value is less than $\lambda$). This gives the flow value. In the former case, Miller and Naor show that the following is a feasible $st$-flow assignment of value $\lambda$. For every edge $d\in G$ that maps to $d^{*}=(g,h)\in G^{*}$: the flow on $d$ is set to $\mathsf{dist}(f,h)-\mathsf{dist}(f,g)+\lambda$ if $d$ is directed towards $t$ in $P$, to $\mathsf{dist}(f,h)-\mathsf{dist}(f,g)-\lambda$ if $d$ is directed towards $s$ in $P$, and to $\mathsf{dist}(f,h)-\mathsf{dist}(f,g)$ if it is not in $P$.
@@
The undirected maximum $st$-flow problem asks to find the maximum flow that can be pushed from $s$ to $t$ in an undirected network. Alternatively, where two neighbors are connected by two opposite direction edges with the same capacity. Although the $s$-to-$t$ flow value equals $t$-to-$s$ flow value, the $s$-to-$t$ flow itself is directed (e.g. the flow entering $s$ is non-positive, the flow entering $t$ is non-negative).
@@
To prove this, we simulate a centralized algorithm of Hassin [20], where he adds an edge $e=(t,s)$ and then computes SSSP in the dual of the resulting graph. $s$ and $t$ are required to be on the same face in order to ensure that $e$ does not violate planarity (it is embedded in the interior of the face containing $s$ and $t$). We use our minor-aggregation simulation on the dual (Section 4) in order to apply an approximate SSSP algorithm of [42] on $(G\cup\{e\})^{*}$, the dual of $G\cup\{e\}$. We use an approximate algorithm since there is no exact one in the minor-aggregation model. This simulation instantly gives the value of the approximate maximum $st$-flow, however, in order to get an assignment we need to work a bit harder because Hassin’s algorithm is an exact algorithm whilst we use it approximately. Hassin defines the flow assignment on an edge $e^{\prime}$ whose dual is $e^{\prime*}=(f,g)$ to be $\mathsf{dist}(f_{1},g)-\mathsf{dist}(f_{1},f)$, where $f_{1}$ is the source of the SSSP tree computed in $(G\cup\{e\})^{*}$ (note, we need the exact distances for the assignment). Thus, approximately, if we want the flow assignment to be capacity respecting (amongst other requirements), for each edge, we need to bound the difference of its endpoints’ (approximate) distances, i.e, we need the approximate distances to obey the triangle inequality. Most (distributed) distance approximation algorithms do not have this feature. Meaning, the approximation can be very far off the true answer. E.g, let $u,v$ be two neighbors of distance $\Omega(n)$ from the source, connected by a zero weight edge, we want the difference of their (approximate) distances from the source to be zero, however, the difference of the approximate distances might be at least as large as $\epsilon n$. To overcome that, we show how to apply a recent algorithm of [40] on the dual. This algorithm was not implemented in the minor-aggregation model before.
@@
The algorithm follows the centralized algorithm of Hassin [20]: Let $f$ be a face of $G$ containing both $s$ and $t$. Firstly, we add an artificial edge $e=(t,s)$ to $G$ embedded inside the face $f$ (preserving planarity) with infinite capacity. This edge is assigned a weight of $nW$, where $W$ is the largest weight in the network, representing infinite capacity in $\tilde{O}(1)$-bits.
@@
Distributively, all vertices need to know the IDs of $s,t$, and $f$. We detect the face $f$ by a simple part-wise aggregation in $\hat{G}$ (Properties 2, 3 and 4 of $\hat{G}$), where the input of each copy of $s$ (resp. $t$) is $ID(s)$ ($ID(t)$) and an identity item for any other vertex (e.g. minus one, as no ID is negative). The aggregate operator simply counts the distinct values of input. Then, the face $f$ is the face that has the minimum ID of the (at most two) faces which got a count of two ($s,t$ know these IDs and can choose the minimum locally). Note, this aggregation terminates in $\tilde{O}(D)$ since the values aggregated are of size $\tilde{O}(1)$-bits and there are at most two distinct values aggregated on each face of $\hat{G}$ that maps to a face of $G$. Then, both $s$ and $t$ add $e$ to the clockwise ordering of their incident edges in $G$, and embed it to be in between two consecutive edges of their incident edges in $f$. We use the minor aggregation model implementation (Section 4) for the dual graph to simulate an SSSP minor-aggregation algorithm on $(G\cup\{e\})^{*}$. We use the algorithm of [42] that computes $(1+\epsilon)$-approximate SSSP in $2^{O(\log n\cdot\log\log n)^{3/4}}/\epsilon^{2}$ minor-aggregation rounds, yielding an $D\cdot n^{o(1)}$ $\mathsf{CONGEST}$ algorithm in $G$ for the undirected $st$ flow problem in $st$-planar networks. However, we want to simulate the algorithm on $(G\cup\{e\})^{*}$ rather than $G^{*}$, meaning the simulation result of Section 4 cannot be used straightforwardly. Moreover, [42]’s algorithm assumes weights in $[1,n^{O(1)}]$ but not zero (not allowing zero capacities). So, we assume that we do not have zero weight edges and first dive into the details of the simulation, then we show how to get an assignment and finally show a simple fix for dealing with zero weight edges.

### v4

@@ 5.1.3 Distributed Knowledge Properties
In $\tilde{O}(D)$ rounds, for each bag $X$, all vertices know for each of their incident darts whether they participate in the critical face (face-part) and whether they participate in any face-part of $X$.
1
Result: Each vertex $v$ of $G$ learns for each incident dart $d$ the list of faces and face-parts that contain $d$ in each bag $X\in\mathcal{T}$ that contains $d$.
2
3Each vertex learns for each incident dart $d$ the list of bags that contains $d$ (Lemma 5.10);
4
5Each vertex learns for each incident dart $d$ the list of faces that contain $d$ in the bag $X=G$ (Lemma 5.11);
6
8
9 /* Global step */
11
13
15
16 /* Local Step */
17  Each vertex of $X$ learns locally for each incident dart $d$ the new ID of each face(-part) that contains $d$ in each child bag of $X^{*}$ that contains $d$ (if any);
18
19 end forall
20
Algorithm 1 Distributed Knowledge of Faces
Proof of Lemma 5.12..
Throughout, we assume that for each $X\in\mathcal{T}$ we already computed a BFS tree $T_{x}$ of $X$ rooted at an arbitrary vertex of $X$. The subscript is dropped when $X$ is not ambiguous. This is done in $\tilde{O}(D)$ rounds using the properties of the BDD. The algorithm described in the proof is given in a high-level pseudocode by Algorithm 1.
First, we apply Lemma 5.10 in order for each vertex to learn for each incident dart the bags that contain it in the decomposition. Then, for $X=G$ we apply Lemma 5.11 assigning each dart a face in $G$. For a general bag $X$, we assume we have the desired information and show how to maintain it for child bags of $X$. For a face or face-part $f$ that is contained in $X$ and is entirely contained in one of its child bags $X_{i}$, endpoints of darts in $f$ inherit the ID of $f$ from $X$. However, in the case where $f$ gets partitioned between child bags of $X$, we need to identify each new face-part and assign it a unique ID.
1. Global step. In this step we detect the face of $G$ and all face-parts $f$ in $X$ to be partitioned (if any). We start with the unique (by Lemma 5.3) critical face (or face-part) of $G$ in $X$. The endpoints of $e_{\!{}_{X}}$ lie on $f$, are on $S_{\!{}_{X}}$ and know if $e_{\!{}_{X}}$ is virtual or not (if it is, then $f$ gets partitioned). Therefore, $f$ can be detected by the endpoints $u,v$ of $e_{\!{}_{X}}$ by the BDD construction (i.e., using the algorithm of [17] that [27] use for computing a separator). More concretely, the edge $e_{\!{}_{X}}$ is embedded between two consecutive incident edges $e_{1},e_{2}$ in the clockwise embedding of $v$ ($e_{1},e_{\!{}_{X}},e_{2}$ appear in that order in the embedding of $v$). Then, the dart of $e_{1}$ that leaves $v$ and the dart of $e_{2}$ that inters $v$ (in that order) define the face $f$. A similar argument holds for $u$. Henceforth, an arbitrary endpoint of $e_{\!{}_{X}}$ broadcasts the ID of the critical face (face-part) on $T$ saying it shall be partitioned as well (if it does).
After identifying the critical face, we move on to identifying the other face-parts that shall be partitioned. We exploit the small (logarithmic) number of face-parts in $X$ (Lemma 5.3) for performing the following broadcast. Each vertex $v$ on a face-part $f$ of $X$ knows (its incident darts in) $f$. Then, $v$ upcasts $ID(f),ID(X_{i})$ to the root of the BFS tree $T$ of $X$, where $X_{i}$ is a child bag that contains darts of $f$ incident to $v$ (the correspondence between darts and child bags is known from Lemma 5.10). We need to be careful in this upcast not to suffer unnecessary congestion, as a face-part might consist of many vertices and more than one connected component. So, if some vertex of $X$ receives the same message more than once, it passes it on only the first time. In particular, if a vertex of $X$ receives more than one child bag ID for the same face-part $f$, it stops passing messages regarding $f$, and upcasts only one special message to the root of $T_{\!{}_{X}}$ saying that $f$ shall be partitioned. The upcast is performed in a pipelined method. Finally, after the root stops receiving messages, it broadcasts the IDs of all face-parts that shall be partitioned.
2. Local step. For each face or a face-part $f$ that is not partitioned, the vertices lying on $f$ consider the ID of $f$ to be the same as from the parent level and remember that it also exists as a whole in the child bag it is contained in. For a face or a face-part $f$ that is partitioned, the vertices $v$ of $f$ know for each incident dart $d$ that belongs to $f$ in which child bag $X_{i}$ $d$ is contained (Lemma 5.10). Thus, $v$ locally learns $(ID(f),ID(X_{i}))$ as the ID of the face-part that contains $d$ in $X_{i}$. We get that all vertices on a face-part of $f$ in some child bag of $X$ know the same ID for that face-part.
We claim that this procedure terminates within $\tilde{O}(D)$ rounds. First, the IDs of face-parts are of size $\tilde{O}(1)$ since (1) by Lemma 5.11 the faces in $G$ are assigned IDs of size $O(\log n)$, (2) by the construction of the BDD in [27], the ID of a child bag is of size $\tilde{O}(1)$ (and there are $O(\log n)$ levels of $\mathcal{T}$), and (3) a face-part in a level $\ell$ is assigned the ID of a face (or face-part) of level $\ell-1$ after appending $\tilde{O}(1)$ additional bits. Second, by Lemma 5.3 the number of face-parts in $X$ is at most $O(\log n)$ and there is at most one critical face (or face-part). Meaning, any vertex shall pass at most $O(\log n)$ messages of size $\tilde{O}(1)$ throughout the course of the procedure (one message for the critical face and one per face-part of $X$). Finally, the BFS tree used for broadcast has a diameter of $O(D\log n)$ (Property 5 of the BDD), and all bags $X$ of the same level run the procedure in parallel with a constant factor overhead (Property 7 of the BDD). ∎
@@
After each vertex $v$ knows the necessary information for each dart of an incident edge $e$, in an additional round, $v$ knows if $e$ exists in $X^{*}$, its weight in $X^{*}$ and its direction, constituting property 14 of the BDD.
@@
Proof.
Recall that by the definition of the dual bag, there is a dual edge $e^{*}$ between two faces or face-parts $f_{1},f_{2}$ of $X$ (on both sides of $e$) iff the two darts of $e$ are contained in $X$ (i.e., none of them is on a hole). By invoking Lemma 5.12 and Lemma 5.10 $v$ knows which of its incident edges $e$ has both darts in $X$, and the ID of the face (face-part) each dart participates in. The endpoints of such edges $e$ compute the correspondence locally as follows. Consider such an edge $e$ incident to a vertex $v\in X$ and let $f_{1},f_{2}$ be the faces containing the darts $d^{+},d^{-}$ of $e$, respectively. Then, the corresponding dual edge is directed from $f_{1}$ to $f_{2}$. Notice, this definition is consistent (independent from $v$) and relies only on the direction of $e$ and on faces it participate in. If $G$ is undirected, then dual edges are undirected as well. If $G$ is weighted, then, $e^{*}$ has the same weight as $e$. ∎
@@ 5.2 The Labeling Scheme
Recall that, for every dual bag $X^{*}$, we wish to label the nodes of $X^{*}$, such that, from the labels alone of any two nodes we can deduce their distance in $X^{*}$. Recall, we refer to the dual node $f^{\prime}$ in a child bag $X^{*}_{i}$ as a node-part of a node $f\in X^{*}$, if $f^{\prime}$ corresponds to a face-part in $X_{i}$ resulting from partitioning the face (face-part) $f$ of $X$.
We already know that the set $F_{\!{}_{X}}$ plays the role of a node-cut (separator) of $X^{*}$, and is crucial for the labeling scheme and algorithm. The main structural property that $F_{\!{}_{X}}$ provides is: Any path in $X^{*}$ either intersects $F_{\!{}_{X}}$ or is entirely contained in a child bag of $X^{*}$.
Note, a node $g\notin F_{\!{}_{X}}$ of $X^{*}$ corresponds to real faces of $G$ (i.e. not face-parts) contained in $X$. The distance label $\mathsf{Label}_{X^{*}}(g)$ of $g\notin F_{\!{}_{X}}$ is defined recursively. If $X^{*}$ is a leaf-bag then $\mathsf{Label}_{X^{*}}(g)$ stores ID($g$), ID($X$) and the distances between $g$ and all other nodes $h\in X^{*}$. Otherwise, the label consists of the ID of $X$, the distances in $X^{*}$ between $g$ and all nodes of $F_{\!{}_{X}}$, and (recursively) the label of $g$ in the unique (by Corollary 5.6) child bag $X^{*}_{i}$ of $X^{*}$ that entirely contains $g$. Formally:
@@
Let $g,h$ be two nodes in a non-leaf bag $X^{*}$, the (edges of the) shortest $g$-to-$h$ path in $X^{*}$ is either entirely contained in some $X^{*}_{i}$ or intersects $F_{\!{}_{X}}$.
Thus, a path in $X^{*}$ from $g\in X^{*}_{i}$ to $h\in X^{*}_{j\neq i}$ must pass through some (part of a) node $f\in F_{\!{}_{X}}$. I.e., a $g$-to-$h$ shortest path $P$ in $X^{*}$ is either: (1) Entirely contained in $X^{*}_{i}$, then $\mathsf{dist}_{X^{*}}(g,h)$ can be deduced from $\mathsf{Label}_{X^{*}_{i}}(g)$ and $\mathsf{Label}_{X^{*}_{i}}(h)$. Or (2), $P$ intersects $F_{\!{}_{X}}$, then $\mathsf{dist}_{X^{*}}(g,h)=\min_{f\in F_{\!{}_{X}}}\{\mathsf{dist}_{X^{*}}(g,f)+\mathsf{dist}_{X^{*}}(f,h)\}$. If one of $g$ or $h$ is a node of $F_{\!{}_{X}}$ then distances are retrieved instantly, as each node in $X^{*}$ stores its distances to $F_{\!{}_{X}}$.
By this reasoning, obtaining the distance between two nodes $g,h$ in a leaf bag is trivial, as all pairwise distances are saved in each label. For a non-leaf bag $X^{*}$, by Lemma 5.15, the information saved in the labels of $g,h$ in $X^{*}$ is sufficient in order to obtain the $g$-to-$h$ distances in $X^{*}$, constituting the following lemma.
@@
The label of any node $g$ in a bag $X^{*}$ of $X\in\mathcal{T}$ is of size $\tilde{O}(D)$ bits.
@@ 5.3 The Labeling Algorithm
3forall bags $X$ of level $\ell\in[\mathsf{Depth}(\mathcal{T}),1]$, in parallel, do
@@ 11     Broadcast the dual edges of $S_{\!{}_{X}}$ edges that exist in $X^{*}$ to the entire bag $X$;
19  Each vertex $v\in X$ computes locally $\mathsf{Label}_{X^{*}}(g)$ for each face (face-part) $g$ of $X$ that contains it or reports a negative cycle aborting the algorithm;
@@
  • An edge between every two nodes of $DDG(g)$ that belong to the same child bag of $X^{*}$. The weight of such an edge $(f,h)\in V_{i}\times V_{i}$ is the $f$-to-$h$ distance in the graph $X^{*}_{i}$. By Lemma 5.16, each such distance can be obtained by pairwise decoding two labels of nodes (received in the broadcast step) that belong to the same child bag $X^{*}_{i}$.
@@
  • A bidirectional zero weighted edge between each two nodes $f_{i}\in V_{i}$, $f_{j}\in V_{j}$ that correspond to node-parts in $X_{i}$, $X_{j}$ (respectively) of the same node $f$ in $F_{\!{}_{X}}$. Again, this can be known locally in any vertex $v$ of $X$ since $v$ received the labels (which contain the IDs) of $f_{i},f_{j}$ in the broadcast step, and since the correspondence of $f_{i},f_{j}$ to $f$ can be known from the information received in the broadcast (specifically, the IDs of $f_{i},f_{j}$ contain the ID of $f$ by Lemma 5.12).
@@
To conclude, a BDD $\mathcal{T}$ is built in $\tilde{O}(D)$ rounds with each vertex $v\in G$ knowing all relevant information that is locally needed about $X^{*}$ (i.e., $X^{*}$ is distributively stored). In each level of the $O(\log n)$ levels of $\mathcal{T}$ we have at most $\tilde{O}(D^{2})$ rounds. Thus, by Lemmas 5.18 and 5.19 the above gives a correct distance labeling algorithm of the dual network. Formally,
@@ 5.4 SSSP in the dual
To mark the SSSP tree edges, an aggregate operator is computed for each node $f$ of $G^{*}$ over its neighbors (using Lemma 4.9, we perform a single part-wise aggregation task). In particular, for a dual node, we mark its incident edge that minimizes the distance product with its neighbors, i.e, we take the operator computed is $\min_{g\in N(f)}\{\mathsf{dist}(s,g)+w(g,f)\}$ (where $N(f)$ is $f$’s neighborhood). Finally, $G^{*}$ might be a multi-graph, so there might be many edges connecting $f$ with a neighbor. We mark the minimal weight edge connecting $f$ to the neighbor that minimizes the above (breaking ties by IDs for edges and for neighbors). Thus,
@@
The above theorem in addition to Section 4 imply distributed (exact and approximate) algorithms for Maximum $st$-Flow in $G$ (i.e., proves Theorems 1.2 and 1.3) by an adaptation of the centralized algorithms of Miller-Naor [31] and of Hassin [20]. The details are given in the next section.
@@ 6.1 Maximum st-Flow
The proof mostly follows a centralized algorithm of Miller and Naor [31], the general idea is to perform a binary search on the value $\lambda$ of the maximum $st$-flow by a logarithmic number of SSSP computations in the dual graph, after which one can deduce the value of the max $st$-flow and a flow assignment.
Proof.
The path $P$ can be found in $O(D)$ rounds using a standard breadth-first search (notice that $P$ is a directed path of darts but does not need to be a directed path of edges in $G$), and the dual SSSP computation is done in $\tilde{O}(D^{2})$ rounds as we showed in Section 5 (with a slight modification that we show next). Since $\lambda$ is assumed to be polynomial, all iterations of the binary search are done in $\tilde{O}(D^{2})$ rounds. The flow assignment on the edges of $G$ is given by the assignment to its darts, that can be deduced locally from its endpoint vertices by Property 14 of the BDD (since for a dart $d=(u,v)\in G$ with $d^{*}=(f,g)\in G^{*}$ the endpoints $u,v$ belong to both faces $f$ and $g$ in $G$).
Notice that the SSSP algorithm on $G^{*}$ from Section 5 uses darts for obtaining the necessary distributed knowledge. However, the distance computations it performs considers edges and not darts. We next show that these computations can be easily modified to work with darts. I.e., to consider the two darts of an edge (with weights as described earlier) when computing the labeling and the shortest paths tree. To do that, we slightly modify the input network and use the algorithm as a black-box on it. That is, we work with the network $G^{\prime}$ obtained from $G$ by adding the reversed dart of each edge to the graph, such that, the two darts of an edge are present in $G^{\prime}$ and its dual. The reversed dart of an edge gets assigned a capacity of zero. It is not hard to see that $G^{\prime}$ can be constructed and embedded in the plane without any need for communication. This is because $G$ is embedded and the two darts $e,\mathsf{rev}(e)$ that map to the edge $e$ are embedded one over the other (the local cyclic ordering of a vertex for its incident edges extends naturally to darts). In addition, each $\mathsf{CONGEST}$ round on $G^{\prime}$ is simulated within two rounds on $G$. Then, the dual $G^{\prime*}$ of $G^{\prime}$ looks exactly like $G^{*}$ after augmenting it with an edge $\mathsf{rev}(e)=(g,f)$ of weight zero (the reversed dart) for each existing edge $e=(f,g)$ in $G^{*}$. Note, some of those edges $\mathsf{rev}(e)$ that we add are parallel edges, meaning, their endpoints may have an edge connecting them with the same direction of $\mathsf{rev}(e)$ (possibly with a different weight), however, this is fine as $G^{*}$ (before the the augmentation) may have parallel edges. I.e., our dual SSSP algorithm knows to handle that. The graph $G^{\prime*}$ can be learnt distributively in the same manner that $G^{*}$ can be learnt, because each dart that we add in the primal is known to its endpoints. Furthermore, the (extended) BDD of $G^{\prime}$ is the same as that of $G$ since $G^{\prime}$ is $G$ with all of its darts present, and since the algorithm for constructing an extended BDD originally considers all darts of $G$. Hence, the dual SSSP algorithm is simulated on $G^{\prime*}$ in the same $\tilde{O}(D^{2})$ round complexity of running it on $G^{*}$.
Then, the maximum $st$-flow algorithm for $G$ runs as follows. First, $G^{\prime}$ is constructed. Then, an $s$-to-$t$ path $P$ in $G^{\prime}$ is found in $O(D)$ rounds via an undirected breadth-first search starting at $s$. Afterwards, $\lambda=1$ units of flow are pushed on $P$ (in the first iteration), and the capacities of $P$’s darts and their reversals are (locally) changed accordingly. Then, we compute an SSSP tree of $G^{\prime*}$ using Lemma 2.2 with the source $f$ being the minimum ID face of $G^{\prime}$ ($f$ is found in $O(D)$ rounds by aggregating on $G^{\prime}$ the IDs of its faces, which are known by Property 13 of the BDD). Following this, each node $v$ in $G^{\prime}$ knows the distance from $f$ to each of the (dual nodes mapped to) faces it lies on. Hence, $v$ learns the flow value for each of its incident darts. Note, in each iteration, each vertex in $G^{\prime}$ knows locally the current value of $\lambda$, depending on whether or not it has received a message reporting a dual negative cycle from the dual SSSP algorithm. When the algorithm termintes, each vertex $v$ of $G$ knows the $st$-flow value in $G$ and (locally) a corresponding flow assignment to its incident edges (the assignment induced by the darts of $G^{\prime}$). ∎
@@
Next we show how to use Section 4 for computing the undirected maximum $st$-flow when $s$ and $t$ lie on the same face of $G$. The undirected maximum $st$-flow problem asks to find the maximum flow that can be pushed from $s$ to $t$ in an undirected network. Alternatively, where each two neighbors are connected by two opposite direction edges with the same capacity. Although the $s$-to-$t$ flow value equals $t$-to-$s$ flow value, the $s$-to-$t$ flow itself is directed (e.g. the flow entering $s$ is non-positive, the flow entering $t$ is non-negative).
@@
To prove this, we simulate a centralized algorithm of Hassin [20], where he adds an edge $e=(t,s)$ and then computes SSSP in the dual of the resulting graph. $s$ and $t$ are required to be on the same face in order to ensure that $e$ does not violate planarity (it is embedded in the interior of the face containing $s$ and $t$). We use our minor-aggregation simulation on the dual (Section 4) in order to apply a $(1+\epsilon)$-approximate SSSP algorithm of [43] on $(G\cup\{e\})^{*}$, the dual of $G\cup\{e\}$. We use an approximate algorithm since there is no exact one in the minor-aggregation model. This simulation instantly gives the value of the approximate maximum $st$-flow, however, in order to get an assignment we need to work a bit harder because, Hassin’s algorithm is an exact algorithm whilst we use it approximately. Hassin defines the flow assignment on an edge $e^{\prime}$ whose dual is $e^{\prime*}=(g,h)$ to be $\mathsf{dist}(f_{1},h)-\mathsf{dist}(f_{1},g)$, where $f_{1}$ is the source of the SSSP tree computed in $(G\cup\{e\})^{*}$ (note, we need the exact distances for the assignment). Thus, approximately, if we want the flow assignment to be capacity respecting (amongst other requirements), for each edge, we need to bound the difference of its endpoints’ (approximate) distances, i.e, we need the approximate distances to obey the triangle inequality. Most (distributed) distance approximation algorithms do not have this feature. Meaning, the approximate flow assignment can be very far off the true answer. E.g, let $u,v$ be two neighbors of distance $\Omega(n)$ from the source, connected by a zero weight edge, we want the difference of their (approximate) distances from the source to be zero, however, the difference of the approximate distances might be at least as large as $\epsilon n$. To overcome that, we show how to apply a recent algorithm of [41] on the dual. This algorithm was not implemented in the minor-aggregation model before.
@@
The algorithm follows the centralized algorithm of Hassin [20]: Let $f$ be a face of $G$ containing both $s$ and $t$. First, we add an artificial edge $e=(t,s)$ to $G$ embedded inside the face $f$ (preserving planarity) with infinite capacity. This edge is assigned a weight of $nW$, where $W$ is the largest weight in the network, representing infinite capacity in $\tilde{O}(1)$-bits.
@@
Distributively, all vertices need to know the IDs of $s,t$, and $f$. We detect the face $f$ by a simple part-wise aggregation in $\hat{G}$ (Properties 2, 3 and 4 of $\hat{G}$), where the input of each copy of $s$ (resp. $t$) is $ID(s)$ ($ID(t)$) and an identity item for any other vertex (e.g. minus one, as no ID is negative). The aggregate operator simply counts the distinct values of input. Then, the face $f$ is the face that has the minimum ID of the (at most two) faces which got a count of two ($s,t$ know these IDs and can choose the minimum locally). Note, this aggregation terminates in $\tilde{O}(D)$ since the values aggregated are of size $\tilde{O}(1)$-bits and there are at most two distinct values aggregated on each face of $\hat{G}$ that maps to a face of $G$. Then, both $s$ and $t$ add $e$ to the clockwise ordering of their incident edges in $G$, and embed it to be in between two consecutive edges of their incident edges in $f$. We use the minor aggregation model implementation (Section 4) for the dual graph to simulate an SSSP minor-aggregation algorithm on $(G\cup\{e\})^{*}$. We use the algorithm of [43] that computes $(1+\epsilon)$-approximate SSSP in $2^{O(\log n\cdot\log\log n)^{3/4}}/\epsilon^{2}$ minor-aggregation rounds, yielding a (1o(1))(1-o(1)) approximate $D\cdot n^{o(1)}$ round $\mathsf{CONGEST}$ algorithm in $G$ for the undirected $st$ flow problem in $st$-planar networks after setting $\epsilon$ to $1/n^{o(1)}$. However, we want to simulate the algorithm on $(G\cup\{e\})^{*}$ rather than $G^{*}$, meaning the simulation result of Section 4 cannot be used straightforwardly. Moreover, [43]’s algorithm assumes weights in $1,n^{O(1)}]$ but not zero (not allowing zero capacities). So, we assume that we do not have zero weight edges and first dive into the details of the simulation, then we show how to get an assignment and finally show a simple fix for dealing with zero weight edges.